OpenStudy (anonymous):

complete order pairs x=8-y using x being 0,6,8 solve for y for each pair

jimthompson5910 (jim_thompson5910):

Simply replace x with 0. Then solve for y x = 8-y 0 = 8-y 0-8 = -y -8 = -y -y = -8 y = 8 So one ordered pair is (0, 8)

jimthompson5910 (jim_thompson5910):

Do the same with x = 6 and x = 8

OpenStudy (anonymous):

i got it i was confused when graphing

jimthompson5910 (jim_thompson5910):

alright tell me what you got for the other two ordered pairs

OpenStudy (anonymous):

i got (6,2) and (8,0)

jimthompson5910 (jim_thompson5910):

you got the correct ordered pairs

jimthompson5910 (jim_thompson5910):

did you clear up the confusion you had when graphing?

OpenStudy (anonymous):

yes

jimthompson5910 (jim_thompson5910):

alright great

OpenStudy (anonymous):

thanks

jimthompson5910 (jim_thompson5910):

you're welcome